OpenStudy (anonymous):

Macy currently has a balance of $3,418.82 in an account earning simple interest. Eighteen years ago she opened the account with an initial deposit of $2,481. What is the interest rate on the account?

OpenStudy (anonymous):

is anyone there any members can help me please

OpenStudy (jdoe0001):

do you happen to have the Interest Rate Formula?

OpenStudy (anonymous):

no :(

OpenStudy (jdoe0001):

or, I should say, if you stick "X" amount in the bank today, and they say they'll give you 5percent interest, how much would you expect in 1year?

OpenStudy (jdoe0001):

would you expect X+[5percent of X]? after 1 year?

OpenStudy (jdoe0001):

ok, what's 10% of 100? :)

OpenStudy (anonymous):

10

OpenStudy (jdoe0001):

ok, so if you put $100 in the bank, and they they'd give you 10 yearly interest, after 1 year, you'd expect $100+[10% of $100 = $10], thus $110 :)

OpenStudy (jdoe0001):

Macy originally put $2481, after 18 years, with yearly interest, it became $3,418.82 so $3,418.82 = $2481 + [$2481 * RATE] *18years

OpenStudy (anonymous):

it ask for the interest rate

OpenStudy (anonymous):

7.7% 1.5% 5.6% 2.1%

OpenStudy (anonymous):

Even if it's earning simple interest, the interest will compound annually.

OpenStudy (jdoe0001):

$$ A= 3,418.82\\ P= 2,481\\ r = ? -------------\\ A=P+ Prt \implies P(1+rt)\\ A-P=Prt\implies \color{blue}{\cfrac{A-P}{Pt}=r} $$

OpenStudy (jdoe0001):

woops, ahemm

OpenStudy (anonymous):

7.7% 1.5% 5.6% 2.1% these are the answrs

OpenStudy (jdoe0001):

$$ A= 3,418.82\\ P= 2,481\\ \color{blue}{r = ?}\\ -------------\\ A=P+ Prt \implies P(1+rt)\\ A-P=Prt\implies \color{blue}{\cfrac{A-P}{Pt}=r} $$

OpenStudy (jdoe0001):

@jcdavi34 what does that give you?

OpenStudy (jdoe0001):

ohh,and T=18, 18years :/

OpenStudy (anonymous):

i got 2.1%

OpenStudy (jdoe0001):

$$ A= 3,418.82\\ P= 2,481\\ t=18\\ \color{blue}{r = ?}\\ -------------\\ A=P+ Prt \implies P(1+rt)\\ A-P=Prt\implies \color{blue}{\cfrac{A-P}{Pt}=r} $$ what do you get?

OpenStudy (jdoe0001):

2.1% that is :)

OpenStudy (anonymous):

thanks!